You are driving in your car and have just purchased a tray of drinks which are sitting on the flat seat next to you. If you are travelling at 15.0 m/s and approach a turn with a radius of 30.0 m, what minimum coefficient of static friction is necessary to keep the drinks from sliding?


1
Expert's answer
2021-02-10T10:09:27-0500
"F_{fr}=F_c,""\mu mg=\dfrac{mv^2}{R},""\mu=\dfrac{v^2}{gR},""\mu=\dfrac{(15.0\ \dfrac{m}{s})^2}{9.8\ \dfrac{m}{s^2}\cdot30\ m}=0.76"